# line of sight CSI channel state information

### v1

@@ I-A The FR3 Spectrum
As the global wireless industry advances towards $6$G, the upper mid-band frequencies, being a part of the high super high frequency (SHF) bands, are critical resources due to their balance between coverage and bandwidth availability, compared to lower frequency range 1 (FR1) (up to $7.125\operatorname{GHz}$ [3]) and higher frequency range 2 (FR2) ranges, which includes $24.25\operatorname{GHz}$ to $71\operatorname{GHz}$ [4]. Often termed the “Golden Band”, or ”Goldilocks Spectrum”, for $6$G, FR3 frequencies ($7.125$-$24.25$ $\operatorname{GHz}$ [2] as depicted in Fig. 1 [1]) are particularly suited for enhancing network capacity while maintaining reasonable propagation characteristics, offering moderate propagation losses, enabling extensive urban and suburban reach using existing towers.
The FR3 spectrum holds growing importance and interest for industry and regulatory bodies like the U.S. National Telecommunications and Information Administration (NTIA), the $3^{\text{rd}}$ Generation Partnership Program (3GPP) [5] and the Federal Communications Commission (FCC) , who are currently evaluating its potential alongside existing mobile radio bands in order to expand cellular services. Besides mobile operator use, the upper mid-band is being considered to coexist with incumbent services such as satellite communications, radio astronomy, and earth exploration, in addition to warfare activities, such as the next generation jammer mid-band (NGJ-MB) operating within $509\operatorname{MHz}$ to $18\operatorname{GHz}$ band [6].
@@ II-A FR3 Channel Modeling
The findings of FR3 channel measurements conducted by NYU WIRELESS in [1, 7, 8] reveal that for wideband channels at $16.95\operatorname{GHz}$ in the FR3 band, the omnidirectional pathloss exponent (PLE) values, synthesized from directional channel measurements, in LoS scenarios were slightly lower than those in milli-meter wave (mmWave), indicating less signal attenuation and more of a waveguide effect over distance, e.g. the omnidirectional LoS PLEs in InH, UMi and InF for $6.75\operatorname{GHz}$, $16.75\operatorname{GHz}$, and $28\operatorname{GHz}$ are given in Table I, e.g. $1.34$ at $6.75\operatorname{GHz}$, $1.32$ at $16.75\operatorname{GHz}$, and $1.2$ at $28\operatorname{GHz}$ [1]. In UMi, the omnidirectional LoS PLEs are $1.79$ at $7\operatorname{GHz}$ [1], $1.85$ at $16.75\operatorname{GHz}$ [1], and $2.02$ at $28\operatorname{GHz}$ [1], which indicates slightly less loss over distance than free space channel (eg when PLE $=2$) and with less loss at lower frequencies. Although antenna patterns and gains vary widely, and higher frequencies enable the use of directional antennas that can offset channel loss in mmWave bands, 3GPP and researchers often use an omnidirectional channel model to standardize link analysis and ensure consistency in evaluations while enabling the use of any type of directional pattern to be applied to the models [10], [11]. NLoS PLEs were found to be higher than LoS scenario as is found in all 3GPP bands as per Table I, e.g. NLoS PLE for UMi of $2.56$ at $6.75\operatorname{GHz}$ [1], $2.59$ at $16.75\operatorname{GHz}$ [1], and for InF the NLoS PLE is $1.78$ at $6.75\operatorname{GHz}$ [8], and $2.11$ at $16.75\operatorname{GHz}$ [8], the loss over distance is much lower (eg $14.5\operatorname{dB}$ per decade less) than those observed at mmWave frequencies, e.g. NLoS PLE of $3.56$ at $28\operatorname{GHz}$, highlighting how FR3 offers improved coverage for a comparable equivalent isotropic radiated power (EIRP) and radio frequency (RF) bandwidth relative to higher frequencies. Referring to Table I, the PLE for LoS in UMi is $1.79$ for $6\operatorname{GHz}$ and $2.02$ for $28\operatorname{GHz}$, implying greater coverage with a $2.2\operatorname{dB}$ stronger signal per decade of distance, resulting in, $4.4\operatorname{dB}$ over $100\operatorname{m}$ and $6.6\operatorname{dB}$ stronger signal over $1\operatorname{km}$, which clearly demonstrates how a lower PLE translates into significant coverage improvement, which may be traded for higher capacity for users within a cell, a factor that will be of paramount importance to carriers when deploying $6$G technologies.
@@
The time delay spread of a propagation channel has impact on signaling formats, in addition to accuracy for integrated sensing and communications (ISAC) applications such as position location and multi-user synchronization methods. Measurements in [1], [7], [8] show that root mean square (RMS) delay spread (DS) decreases with carrier frequency, with FR3 exhibiting smaller DS values than sub-$6$ $\operatorname{GHz}$, but larger than at FR2 mmWave frequencies, e.g. $14\operatorname{ns}$ and $12.7\operatorname{ns}$ LoS omnidirectional RMS DS at $6.75\operatorname{GHz}$ and $16.95\operatorname{GHz}$, respectively for InF. The range of RMS DS was found to be $12.7-37.7\operatorname{ns}$ across InF, UMi and InH. DS values observed in the field indicates that multipath components are closer in time at FR3, suggesting limited temporal dispersion which could favor high-speed, low-latency communications and more accurate position location and timing methods in dense environments. Furthermore, the measurements in [1] found that higher FR3 frequencies, i.e. $16.75\operatorname{GHz}$, exhibited a narrower RMS angular spread compared to the higher FR1 frequency at $6.75\operatorname{GHz}$, hence indicating fewer, more focused multipath components, which is beneficial for spatial multiplexing as it reduces interference between signal paths and allows more precise beamforming. Regarding material penetration, losses were consistently higher at $16.95\operatorname{GHz}$ in FR3 than at lower frequencies, with losses dependent on material type and polarization configuration, confirming the inherent limitations of FR3 in penetrating certain materials, such as low-emissivity (IRR) glass and concrete [1], yet able to allow more penetration that at FR3 mmWave [1].
@@ III-A Spectrum Agility for Coverage-Rate Tradeoffs
The lower FR3 frequencies provide better coverage, as detailed in Section II, hence a fundamental question arises: What mechanism can optimally balance rate-coverage tradeoffs ? A particularly interesting ray tracing simulation was conducted in [14] to study the coverage and rate statistical behaviour across the FR3 spectrum. In terms of rate, with no blockages, and as expected, it was shown that $24\operatorname{GHz}$ usage dominates in terms of rate for outdoors. Interestingly, when a 3GPP blockage model B was used, it was shown that there was a fallback need towards $12\operatorname{GHz}$ and $6\operatorname{GHz}$ for maximum rate. But even more, and making use of the the comprehensive and detailed penetration models and findings from NYU WIRELESS [1], it is observed that while materials like glass and wood exhibit relatively permeable characteristics at FR3 frequencies, others, e.g. low-emissivity glass, are not permeable. A particularly intriguing observation is that concrete, which is relatively permeable at sub-$6$ $\operatorname{GHz}$ frequencies, becomes completely impermeable at FR3. Additionally, as reported in [14], the indoor user rate distribution indicates that users operating at $24\operatorname{GHz}$ will need to fallback to lower frequencies, which means that, if we really want to leverage the large FR3 spectrum, frequency hopping mechanisms between FR3 bands should be designed to account for the various blockages and penetration losses. The diversity and hopping mechanism naturally pushes forward the need for a spectrum agility in the FR3. Besides, the agility also inherently secures certain transmissions from malicious users and eavesdroppers. A natural question then to ask would be what hopping mechanism should be adopted between the FR3 bands in order to be spectrally efficient and agile?

### v2

@@ I-A The FR3 Spectrum
As the global wireless industry advances towards $6$G, the upper mid-band frequencies, being a part of the high super high frequency (SHF) bands, are critical resources due to their balance between coverage and bandwidth availability, compared to lower frequency range 1 (FR1) (up to $7.125\operatorname{GHz}$) and higher frequency range 2 (FR2) ranges, which includes $24.25\operatorname{GHz}$ to $71\operatorname{GHz}$ [3]. Often termed the “Golden Band”, or ”Goldilocks Spectrum”, for $6$G, FR3 frequencies ($7.125$-$24.25$ $\operatorname{GHz}$ [2] as depicted in Fig. 1 [1]) are particularly suited for enhancing network capacity while maintaining reasonable propagation characteristics, offering moderate propagation losses, enabling extensive urban and suburban reach using existing towers.
The FR3 spectrum holds growing importance and interest for industry and regulatory bodies like the U.S. National Telecommunications and Information Administration (NTIA), the $3^{\text{rd}}$ Generation Partnership Program (3GPP) and the Federal Communications Commission (FCC) , who are currently evaluating its potential alongside existing mobile radio bands in order to expand cellular services. Besides mobile operator use, the upper mid-band is being considered to coexist with incumbent services such as satellite communications, radio astronomy, and earth exploration, in addition to warfare activities, such as the next generation jammer mid-band (NGJ-MB) operating within $509\operatorname{MHz}$ to $18\operatorname{GHz}$ band.
@@ II-A FR3 Channel Modeling
The findings of FR3 channel measurements conducted by NYU WIRELESS in [1, 4, 5] reveal that for wideband channels at $16.95\operatorname{GHz}$ in the FR3 band, the omnidirectional pathloss exponent (PLE) values, synthesized from directional channel measurements, in LoS scenarios were slightly lower than those in milli-meter wave (mmWave), indicating less signal attenuation and more of a waveguide effect over distance, e.g. the omnidirectional LoS PLEs in InH, UMi and InF for $6.75\operatorname{GHz}$, $16.95\operatorname{GHz}$, and $28\operatorname{GHz}$ are given in Table I, e.g. $1.34$ at $6.75\operatorname{GHz}$, $1.32$ at $16.95\operatorname{GHz}$, and $1.2$ at $28\operatorname{GHz}$ [1]. In UMi, the omnidirectional LoS PLEs are $1.79$ at $6.75\operatorname{GHz}$ [1], $1.85$ at $16.95\operatorname{GHz}$ [1], and $2.02$ at $28\operatorname{GHz}$[7], which indicates slightly less loss over distance than free space channel (e.g., when PLE $=2$) and with less loss at lower frequencies. Although antenna patterns and gains vary widely, and higher frequencies enable the use of directional antennas that can offset channel loss in mmWave bands, 3GPP and researchers often use an omnidirectional channel model to standardize link analysis and ensure consistency in evaluations while enabling the use of any type of directional pattern to be applied to the models [9], [10]. NLoS PLEs were found to be higher than LoS scenario as is found in all 3GPP bands as per Table I, e.g. NLoS PLE for UMi of $2.56$ at $6.75\operatorname{GHz}$ [1], $2.59$ at $16.95\operatorname{GHz}$ [1], and for InF the NLoS PLE is $1.78$ at $6.75\operatorname{GHz}$ [5], and $2.11$ at $16.95\operatorname{GHz}$ [5], the loss over distance in NLoS locations is lower (e.g., $12$ to $14.5\operatorname{dB}$ per decade less) than those observed at mmWave frequencies, e.g. NLoS PLE of $3.4$ to $3.56$ at $28\operatorname{GHz}$ UMi [7], highlighting how FR3 offers improved coverage in NLOS channels relative to higher-frequency mmWave bands. It is important to note, however, that practical systems use higher antenna gains at higher frequencies, and a common error is to attempt to use equivalent isotropic radiated power (EIRP) to compare coverage at different frequencies, since the transmit and receive antenna physical apertures are proportionally reduced with shrinking wavelength if antenna gains are left constant over frequency, such that equal power transmitters with EIRP over different frequency will give the false impression that higher bands have far less coverage. In practice, the proper comparison is to consider equal antenna aperture areas at different frequencies, where in practice the extra array gain at mmWave can overcome the additional channel loss by many dB [11].
Referring to Table I, the PLE for LoS in UMi is $1.79$ for $6.75\operatorname{GHz}$ and $2.02$ to $2.1$ [7] for $28\operatorname{GHz}$, implying greater coverage with a $2.2\operatorname{dB}$ stronger signal per decade of distance, resulting in, $4.4\operatorname{dB}$ over $100\operatorname{m}$ and $6.6\operatorname{dB}$ stronger signal over $1\operatorname{km}$, which clearly demonstrates how a lower PLE translates into significant coverage improvement, which may be traded for higher capacity for users within a cell, a factor that will be of paramount importance to carriers when deploying $6$G technologies.
@@
The time delay spread of a propagation channel has impact on signaling formats, in addition to accuracy for integrated sensing and communications (ISAC) applications such as position location and multi-user synchronization methods. Measurements in [1], [4], [5] show that root mean square (RMS) delay spread (DS) decreases with carrier frequency, with FR3 exhibiting smaller DS values than sub-$6$ $\operatorname{GHz}$, but larger than at FR2 mmWave frequencies, e.g. $14\operatorname{ns}$ and $12.7\operatorname{ns}$ LoS omnidirectional RMS DS at $6.75\operatorname{GHz}$ and $16.95\operatorname{GHz}$, respectively for InF. The range of RMS DS was found to be $12.7-37.7\operatorname{ns}$ across InF, UMi and InH. DS values observed in the field indicates that multipath components are closer in time at FR3, suggesting limited temporal dispersion which could favor high-speed, low-latency communications and more accurate position location and timing methods in dense environments. Furthermore, the measurements in [1] found that higher FR3 frequencies, i.e. $16.95\operatorname{GHz}$, exhibited a narrower RMS angular spread compared to the higher FR1 frequency at $6.75\operatorname{GHz}$, hence indicating fewer, more focused multipath components, which is beneficial for spatial multiplexing as it reduces interference between signal paths and allows more precise beamforming. Regarding material penetration, losses were consistently higher at $16.95\operatorname{GHz}$ in FR3 than at lower frequencies, with losses dependent on material type and polarization configuration, confirming the inherent limitations of FR3 in penetrating certain materials, such as low-emissivity (IRR) glass and concrete [1], yet able to allow more penetration that at FR3 mmWave [1].
@@ III-A Spectrum Agility for Coverage-Rate Tradeoffs
Under a constant EIRP with shrinking aperture assumption, the lower FR3 frequencies provide better coverage, as detailed in Section II. If instead the antenna physical area is held fixed, the extra array gain available at mmWave more than compensates for the additional free space pathloss. A fundamental question arises: What mechanism can optimally balance rate-coverage tradeoffs ? A particularly interesting ray tracing simulation was conducted in [14] to study the coverage and rate statistical behaviour across the FR3 spectrum. In terms of rate, with no blockages, and as expected, it was shown that $24\operatorname{GHz}$ usage dominates in terms of rate for outdoors. Interestingly, when a 3GPP blockage model B was used, it was shown that there was a fallback need towards $12\operatorname{GHz}$ and $6\operatorname{GHz}$ for maximum rate. But even more, and making use of the the comprehensive and detailed penetration models and findings from NYU WIRELESS [1], it is observed that while materials like glass and wood exhibit relatively permeable characteristics at FR3 frequencies, others, e.g. low-emissivity glass, are not permeable. A particularly intriguing observation is that concrete, which is relatively permeable at sub-$6$ $\operatorname{GHz}$ frequencies, becomes completely impermeable at FR3. Additionally, as reported in [14], the indoor user rate distribution indicates that users operating at $24\operatorname{GHz}$ will need to fallback to lower frequencies, which means that, if we really want to leverage the large FR3 spectrum, frequency hopping mechanisms between FR3 bands should be designed to account for the various blockages and penetration losses. The diversity and hopping mechanism naturally pushes forward the need for a spectrum agility in the FR3. Besides, the agility also inherently secures certain transmissions from malicious users and eavesdroppers. A natural question then to ask would be what hopping mechanism should be adopted between the FR3 bands in order to be spectrally efficient and agile?
@@ IV-D Coherence time across FR3
IV-E RF frontend efficiency and array integration limits
Energy efficiency of power amplifiers can becoem a challenge in the FR3 band, so transmit side losses can limit the link budget advantages. Particularly, packing large numbers of elements into a fixed aperture soon hits practical barriers due to thermal dissipation, EIRP limits, and fronthaul capacity. Therefore, from array design and RF frontend perspective, how can future FR3 radios achieve high energy efficiency in their power amplifiers while packing dense antenna arrays under real world thermal, EIRP, and processing limits, without giving up the very coverage-rate gains the band is meant to provide? Waste factor can serve as a fundamental metric towards the design.